# CachePrune: Neural-Based Attribution Defense Against Indirect Prompt Injection Attacks

### v1

@@ 2.4 The Triggering Effect
Feature attribution with $\mathcal{L}^{attr}_{full}$ in (9) relies on the probabilities for all the tokens in $y_{x}^{p,*}$ and $y_{x}^{c,*}$. However, we show that not all the response tokens are necessary for feature attribution. Specifically, we find that the same input context can be treated as data or instruction by LLMs, depending on the generation of only few trigger tokens, e.g., Figure 7, that precede the model response. We call this the triggering effect.
To illustrate, we plot in Figure 3 the prediction rank of tokens from $y_{x}^{p/c,*}$ in the LLM’s prediction. Formally, the rank for token $y_{x,t}^{p/c,*}$ is,
@@
Motivated by Figure 3, we only perform feature attribution with the first $k$ tokens in the response, which has been enough to make the difference between the clean and poisoned responses. We define the final attribution loss function as,
$\begin{split}\mathcal{L}^{attr}=\mathbb{E}_{x\sim\mathcal{X}}\,\,(\,\,p_{\theta}(y^{p,*}_{x,<k+1}|x)-p_{\theta}(y^{c,*}_{x,<k+1}|x)\,\,)\\ \end{split}$ (15)
where we default with $k=1$. In experiments, we show that the expectation term in (15) can be estimated with only $N=8$ samples. In Figure 3, we use $y_{x}^{p,*}$ that starts with answering the injected instruction. Thus, in computing (15) for feature attribution, we construct such $y_{x}^{p,*}$ by adding "Answer this at the end." before the user query. Note that we do not assume our testing data contains such instruction.
@@ 3 Related Works
Indirect Prompt Injection Attack Different from the direct prompt injection attack Perez and Ribeiro (2022); Yu et al. (2023) which straightforwardly inserts undesirable content into the LLM’s prompt, the indirect prompt injection attack occurs when the input context is injected with third-party instructions Liu et al. (2023); Zhan et al. (2024); Wu et al. (2024); Liu et al. (2024). These instructions can be malicious or not, but not intended to be responded to by the LLM. The success of indirect prompt inject exploits the LLM’s inability to distinguish between the data and instruction Greshake et al. (2023), i.e., it happens when the LLM fails to leverage the context as pure data but responding to its instructions.
Defending Against Prompt Injection Attack Previous defenses against prompt injection attacks can be categorized into training-based and testing-based. For the training-based, the LLM that is identified as subject to indirect prompt injection attack will be trained with extra SFT Chen et al. (2024a) or preference data Chen et al. (2024b) that inform the model on input prompt structure over context vs. instructions. For testing based, existing approaches either modify on the original prompt with prompt engineering Wu et al. (2023); Hines et al. (2024), or design complex workflows Wang et al. (2024); Jia et al. (2024) that introduce extra computations or LLM calls. In this paper, we mitigate the attack with a focus on the foundational problem of the discretion between data and instructions. Our CachePrune is compatible with the existing approaches, while not modifying the prompt or introducing extra two time LLM calls.
@@ 4 Experiment
ASR F1(clean) F1 (attack) ASR F1(clean) F1 (attack) ASR GPT-Score
@@
Encode__\__Base64 6.56 13.34 11.56 3.05 4.24 3.19 5.50 1.52
CachePrune 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84$\pm$ 0.18 15.23 $\pm$ 1.56 16.21 $\pm$ 0.61 10.97 $\pm$ 0.35 2.00 $\pm$ 0.41 3.32 $\pm$ 0.10
@@
Table 1: Results of defending against indirect prompt injection attack. Our CachePrune is implemented on Vanilla. The Bold font denotes the best value for each metric. We use italics instead for Encode__\__Base64, since its ASR is at the expense of very bad response quality (very low F1).
@@ 4.1 Experiment Setup
Model and Dataset We evaluate our approach on the model of LLama3-8B Touvron et al. (2023) and Mistral-7B-Instruct-V3.0 Jiang et al. (2023). We by default experiment with $N=8$ for feature attribution and prune with $p=0.5$ ($0.5\%$ neurons). We evaluate with the question answering datasets of SQuAD Rajpurkar (2016) and HotpotQA Yang et al. (2018). We test on the splits of SQuAD and HotpotQA that are directly processed by Abdelnabi et al. (2024), which randomly injects instructions into the beginning, middle, and ending of the context of each prompt. Our approach focuses on the LLM’s fundamental ability to distinguish between data and instructions, making it applicable to problems beyond defense against third-party injections. Specifically, we also explore a practical scenario of dialogue summarization with the WildChat Zhao et al. (2024) dataset. For this task, the model is attacked if it answers the question raised by users in the dialogue, instead of summarizing the dialogue interactions. We use the same split as in Abdelnabi et al. (2024). We find that the models are rarely attacked with plain dialogues. To increase the difficulty, we insert "You should primarily focus on this question" as part of the user instruction to the AI assistant that appeared in the dialogue. For each dataset, we randomly select 8 samples from a pool of  400 prompts that are not overlapped with the testing data. Results are averages with 3 trials.
Metrics We evaluate SQuAD and HotpotQA with the three metrics. Attack Success Rate (ASR) $\downarrow$: The proportion of poisoned responses from greedy decoding. F1 (clean) $\uparrow$: The F1 score without injected instructions. F1 (Attack) $\uparrow$: The F1 score with injected instructions. For the task of dialogue summarization, we replace the F1 scores with an LLM Judge Zheng et al. (2024) that evaluates the quality of generated summaries into scores ranging [1,5]. We call it the GPT-score.
@@
Figure 4: Performance of LLama3-8 on SQuAD with different percentage of pruned neurons $p$.
Baselines. We primarily compare with the following baselines from Wu et al. (2023); Hines et al. (2024). Vanilla: Original prompt without any defense technique. Delimiting: Adding special characters at the start and end of the context. Datamarking: Replace every space in the context with a special character. Encode__\__Base64: The context is encoded into Base64 while the other text spans are provided with plain text. For fair comparison, we do not compare with baselines of finetuning or requiring test-time computation with extra LLM calls per response.
Our CachePrune is implemented based on Vanilla. We should note that our CachePrune is actually complementary to the other baselines, since our approach does not modify the prompt.
(a) LLama3-8B
(b) Mistral-7B
ASR F1 (clean) F1 (attack)
k=1 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
k=2 5.57 $\pm$ 0.30 26.03 $\pm$ 0.28 22.47 $\pm$ 0.37
k=4 10.77 $\pm$ 0.45 24.71 $\pm$ 0.37 19.29 $\pm$ 0.33
$\mathcal{L}^{attr}_{full}$ 14.81 $\pm$ 0.59 25.63 $\pm$ 0.39 19.78 $\pm$ 0.55
ASR F1 (clean) F1 (attack)
$\alpha$=1.5 6.40 $\pm$ 0.32 26.71 $\pm$ 0.53 20.22 $\pm$ 0.56
$\alpha$=1.0 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
$\alpha$=0.5 10.77 $\pm$ 0.61 28.33 $\pm$ 0.40 21.29 $\pm$ 0.61
$\alpha$=0.3 13.50 $\pm$ 0.70 28.91 $\pm$ 0.37 21.78 $\pm$ 0.43
ASR F1 (clean) F1 (attack)
@@
Table 4: Transferring the mask from feature attribution between code-based and text-based attack.
4.2 Result Analysis
Specifically, the ASR with our proposed CachePrune can be several times lower than Vanilla, Delimiting, and Datamarking. The Encode__\__Base64 yields ASR that is comparable to CachePrune, but at the expense of very low F1 scores. We reckon that this is because the modification on context with Encode__\__Base64 is too complex for our LLMs, resulting in the model understanding the context. This highlights a deficiency of defending with prompt engineering, i.e., the manually designed complex marking on the input context may increase the difficulty for the LLM to comprehend the context information. On the contrary, our approach leverage the LLMs’ perspective on the difference between data and instruction, instead of relying on complex human engineering. Additionally, we can find that the score of F1 (attack) is generally lower than F1 (clean), suggesting that responding to the injected instructions could limit the LLMs’ ability to solve the user-specified ones.
In Figure 5, we plot the distribution of pruned neurons across layers. It can be observed that the neurons that are indicative of the data vs instruction concentrate in the middle layers of the LLM. This is aligned with previous studies, e.g., Huang et al. (2024), showing that the middle layers are more capable of capturing abstract and complex concepts. Additionally, it is interesting to find that there are more key neurons being pruned in layers of LLama3-8b, indicating that the LLama model is trained to perceive instructions with the key vectors. In comparison, the Mistral model is more balanced with keys and values in distinguishing between the concepts of data vs. instructions. In Figure 5, we plot the model performance with the prune ratio $p$. It can be observed that the pruning not necessarily decrease the F1 (clean). This could because the masking on context improves the LLM’s undersdanding of the prompt structure (context vs instruction).
In Table 2, we list the performance of LLama3-8B on SQuAD, with different values of $k$. It suggests that the earlier tokens in the generation of the response are more indicative of the model’s decision on data vs. instruction. Table 3 shows the performance with the ratio of masking $\alpha$. This shows that our identified neurons are indeed reflects the model’s own definition of data vs. instruction, with lower ASR corresponding to larger degree of masking ($\alpha\uparrow$). In Table 4, we show with SQUaD that the mask learn from text/code-based attack can be effectively transferred to defend code/text-based attack. We inject on SQuAD context with code-based attack from Chaudhary (2023) and text-based attack from Ji et al. (2023).
@@ 5 Discussion
We presented a lightweight and efficient approach to mitigate the indirect prompt injection attack. By identifying and neutralizing task-triggering neurons in the key-value (KV) cache, our approach enforces the model treat the input context as solely supportive data. Experimental results demonstrate that CachePrune significantly reduces the ASR without compromising output quality.
@@
Appendix A Proof of Theorem 1

### v2

@@ 2.4 The Triggering Effect
Neural attribution with $\mathcal{L}^{attr}_{full}$ in (8) relies on the probabilities for all the tokens in $y_{x}^{p,*}$ and $y_{x}^{c,*}$. However, we show that not all the response tokens are necessary for neural attribution. Specifically, we find that the same input context can be treated as data or instruction by LLMs, depending on the generation of only few trigger tokens, e.g., Figure 8, that precede the model response. We refer to this as the triggering effect.
To illustrate, we plot in Figure 3 the prediction rank of tokens from $y_{x}^{p,*}$ and $y_{x}^{c,*}$ in the LLM’s prediction. For example, the rank for token $y_{x,t}^{p,*}$ is,
@@
Motivated by Figure 3, we only perform neural attribution with the first $k$ tokens in the response, which has been enough to make the difference between the clean and poisoned responses. We define the final attribution loss function as,
[…]
where we default with $k=1$. In experiments, we show that the expectation term in (13) can be effectively estimated with only $N=8$ samples. In Figure 3, we use $y_{x}^{p,*}$ that starts with answering the injected instruction. Thus, in computing (13) for feature attribution, we construct such $y_{x}^{p,*}$ by adding "Answer this at the end." before the user query. Note that we do not assume our testing data contains such instruction.
@@ 3 Related Works
Indirect Prompt Injection Attack Different from the direct prompt injection attack perez2022ignore; yu2023assessing which straightforwardly inserts undesirable content into the LLM’s prompt, the indirect prompt injection attack occurs when the input context is injected with third-party instructions liu2023prompt; zhan2024injecagent; wu2024wipi; liu2024automatic. These instructions can be malicious or not, but not intended to be responded to by the LLM. The success of indirect prompt inject exploits the LLM’s inability to distinguish between the data and instruction greshake2023not, i.e., it happens when the LLM fails to leverage the context as pure data but responding to its instructions.
Defending Against Prompt Injection Attack
There exists prior defense works following a Detection + Filtering approach getmydrift; piguard; embeddingdetect that build classifiers to detect unauthorized injections and discard or refuse to answer such inputs. These methods are orthogonal to ours, as they focus on detection accuracy. Our setup is measured by Attack Success Rate (ASR) and response quality, i.e., requiring the model to still produce a response regardless of whether an attack is present. Within this line of research, existing approaches can be broadly categorized into training-based and testing-based methods. For the training-based, the LLM that is identified as subject to indirect prompt injection attack will be trained with extra SFT chen2024struq or preference data chen2024aligning that inform the model on input prompt structure over context vs. instructions. For testing based, existing approaches either modify on the original prompt with prompt engineering wu2023defending; hines2024defending, or design complex workflows wang2024fath; jia2024task that introduce extra computations or LLM calls. In this paper, we mitigate the attack with a focus on the foundational problem of the discretion between data and instructions. Specifically, we identify neurons that contribute to the difference between data and instructions via neural attribution. Our CachePrune is compatible with the existing approaches, while not modifying the prompt or introducing extra test time LLM calls. In addition, there are previous works (e.g., attribution patching) syed2023attribution; wu2024mitigating studying how to accurately attribute the outputs to the LLM’s hidden states. We want to emphasize that our study aims at leveraging neuron attribution to understand the model’s cognition on data vs. instruction, instead of proposing a completely new neuron intervention technique.
In addition, our masking requires pre-knowing the context boundary. As in piet2024jatmo, this aligns with LLM-integrated APIs in which user queries are combined with third-party data using API-specific templates, i.e., the position of the context span is also known. In addition, this is common with prior works, e.g., chen2024struq; chen2024aligning; piet2024jatmo; abdelnabi2024you, which operates under fixed templates with known context.
@@ 4 Experiment
ASR $\downarrow$ F1(clean) $\uparrow$ F1 (attack) $\uparrow$ ASR $\downarrow$ F1(clean) $\uparrow$ F1 (attack) $\uparrow$ ASR $\downarrow$ GPT-Score $\uparrow$
@@
Sandwich 21.43 27.69 18.98 67.21 15.30 3.99 13.01 3.22
Encode_\_Base64 6.56 13.34 11.56 3.05 4.24 3.19 5.50 1.52
CachePrune 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49 15.23 $\pm$ 1.56 16.21 $\pm$ 0.61 10.97 $\pm$ 0.35 2.00 $\pm$ 0.41 3.32 $\pm$ 0.10
@@
Phi-3.5-mini- instruct (3.8B) Vanilla 10.22 26.03 25.64 21.67 14.14 7.69 3.32 3.78
Delimiting 7.87 26.05 25.49 11.36 13.68 11.11 3.20 3.84
Datamarking 3.54 26.71 26.47 3.24 12.74 9.78 2.53 3.71
Sandwich 18.65 23.78 23.13 40.17 12.56 5.17 4.37 3.56
Encode_\_Base64 0.86 7.87 4.52 0.07 8.42 7.01 3.56 1.09
CachePrune 0.71 $\pm$ 0.18 26.76 $\pm$ 0.56 25.55 $\pm$ 0.60 1.76 $\pm$ 0.50 14.17 $\pm$ 0.78 9.79 $\pm$ 1.12 1.89 $\pm$ 0.25 3.60 $\pm$ 0.31
Table 1: Results of defending against indirect prompt injection attack. Our CachePrune is implemented on Vanilla. Bold font denotes the best value for each metric. We use underscore instead when Encode_\_Base64 has the lowest ASR, since its ASR is at the expense of very bad response quality (very low F1). $\downarrow$ and $\uparrow$ indicate that lower and higher scores are better, respectively. We also experiment with an adaptuve attack in Appendix G.
@@ 4.1 Experiment Setup
Model and Dataset We evaluate our approach on the model of LLama3-8B touvron2023llama, Mistral-7B-Instruct-V3.0 jiang2023mistral and Phi-3.5-mini- instruct (3.8B) abdin2024phi3technicalreporthighly. We by default experiment with $N=8$ for neural attribution and prune with $p=0.5$ ($0.5\%$ neurons). We evaluate with the question answering datasets of SQuAD rajpurkar2016squad and HotpotQA yang2018hotpotqa. We test on the splits of SQuAD and HotpotQA that are directly processed by abdelnabi2024you, which randomly injects instructions into the beginning, middle, and ending of the context of each prompt. Our approach focuses on the LLM’s fundamental ability to distinguish between data and instructions, making it applicable to problems beyond defense against third-party injections. Specifically, we also explore a practical scenario of dialogue summarization with the WildChat zhao2024wildchat dataset. For this task, the model is attacked if it answers the question raised by users in the dialogue, instead of summarizing the dialogue interactions. We use the same split as in abdelnabi2024you. We find that the models are rarely attacked with plain dialogues. To increase the difficulty, we insert "You should primarily focus on this question" as part of the user instruction to the AI assistant that appeared in the dialogue. For each dataset, we randomly select 8 samples from a pool of  400 prompts that are not overlapped with the testing data. Results are averages with 3 trials. Please refer to Appendix A for more details on baseline ans metrics. We should note that our CachePrune is actually complementary to the existing baselines, since our approach does not modify the prompt or requiring additional test time workflows.
@@
Figure 4: Performance of LLama3-8 on SQuAD with different percentage of pruned neurons pp.
(a) ASR
(b) F1 (attack)
(c) f1 (clean)
Figure 5: Transferring the learnt mask between prompts with different attacks instructions on SQUaD with LLama3-8b. 0-3 corrsponds to 4 different attacks instructions described in Appendix F. Each value in the matrix can be formatted with $[Fi,Tj:s]$. $i,j$ means using the mask learn from prompts attacted by $i$ to pruning KV cache from prompts attacked by $j$. $s$ is the vanilla baseline on $j$ without any defense. The value of $[Fi,Tj:s]$ means the gain in ASR or F1 compared to $s$. For example, $[F0,T3:27.86]$ in (a) means the ASR of applying the mask learnt from attack 0 to attack $1$ is 27.86+(16.32)=11.5427.86+(-16.32)=11.54. We have dark color highlight more negative gain for ASR and more positive gain for F1. Therefore, if the learnt masks are not transferrable between different attacks, we would expect three diagonal matrix in color. However, we can see that the darkest color does not necessrily appear in the diagonal. Therefore, our learn mask for pruning is transferrable between different attacks.
4.2 Result Analysis
Specifically, the ASR with our proposed CachePrune can be several times lower than Vanilla, Delimiting, and Datamarking. The Encode_\_Base64 yields ASR that is comparable to CachePrune, but at the expense of very low F1 scores. We reckon that this is because the modification on context with Encode_\_Base64 is too complex for our LLMs, resulting in the model understanding the context. This highlights a deficiency of defending with prompt engineering, i.e., the manually designed complex marking on the input context may increase the difficulty for the LLM to comprehend the context information. On the contrary, our approach leverage the LLMs’ perspective on the difference between data and instruction, instead of relying on complex human engineering. Additionally, we can find that the score of F1 (attack) is generally lower than F1 (clean), suggesting that responding to the injected instructions could limit the LLMs’ ability to solve the user-specified ones.
In Figure 6, we plot the distribution of pruned neurons across layers. It can be observed that the neurons that are indicative of the data vs instruction concentrate in the middle layers of the LLM. This is aligned with previous studies, e.g., huang2024learn, showing that the middle layers are more capable of capturing abstract and complex concepts. Additionally, it is interesting to find that there are more key neurons being pruned in layers of LLama3-8b, indicating that the LLama model is trained to perceive instructions with the key vectors. In comparison, the Mistral model is more balanced with keys and values in distinguishing between the concepts of data vs. instructions. In Figure 6, we plot the model performance with the prune ratio $p$. It can be observed that the pruning not necessarily decrease the F1 (clean). This could because the masking on context inform the LLM of the prompt structure (context vs instruction).
In Table 3, we list the performance of LLama3-8B on SQuAD, with different values of $k$. It suggests that the earlier tokens in the generation of the response are more indicative of the model’s decision on data vs. instruction. Table 4 shows the performance with the different masking parameetr $\alpha$. With $\alpha$ gets larger, the model is more and more treating the context as instructions which is demonstrated by larger ASR. It shows that our identified neurons are indeed reflecting the model’s own definition of data vs. instruction. In Table 2, we show with SQUaD that the mask learn from text/code-based injection can be effectively transferred to defend code/text-based injection. We inject on SQuAD context with code-based injection task from codealpaca and text-based injection task from ji2023beavertails. In Figure 5, we also show that our learnt mask is transferrable between different attack instructions. The detailed config is described in Appendix F.
ASR $\downarrow$ F1 (clean) $\uparrow$ F1 (attack) $\uparrow$
@@ 4.2 Result Analysis
Table 2: Transferring the mask from feature attribution between code-based and text-based injection. Learning a mask from prompt with code/text injections and apply on data with text/code injections.
@@ 5 Discussion
We presented a lightweight and efficient approach to mitigate the indirect prompt injection attack. By identifying and neutralizing task-triggering neurons in the key-value (KV) cache, our approach enforces the model treat the input context as solely supportive data. Experiments show that CachePrune exhibites robustness and generalizability across various attack types, significantly reducing the ASR without compromising output quality. Our work highlights a practical and scalable solution for enhancing the reliability of LLMs in security-critical applications.
@@
6 Limitations
Our work does not explore alternative training-based defense approaches, such as adversarial fine-tuning, since we target the scenario without a heavy computation budget. Though complementary to our work, future work could benefit from a comparative study of test-time versus training-based defenses to better understand the trade-offs between computation and model resilience.
Appendix A Additional Details
Metrics and Evaluation We evaluate SQuAD and HotpotQA with the three metrics. Attack Success Rate (ASR) $\downarrow$: The proportion of poisoned responses from greedy decoding. F1 (clean) $\uparrow$: The F1 score without injected instructions. F1 (Attack) $\uparrow$: The F1 score with injected instructions. For the task of dialogue summarization, we replace the F1 scores with an LLM Judge zheng2024judging that evaluates the quality of generated summaries into scores ranging [1,5], which we denote as the GPT-score. For each dataset, we randomly select $N=8$ samples from a pool of 400 prompts that are not overlapped with the testing data. Those 400 prompts ara also randonly sampled.
Listing 1 shows the code snippet the computes the F1 score in the main paper, following standardized evaluation for SQuAD and HotpotQA. Under the hood, it is computed from:
$\mathrm{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\quad\mathrm{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{FN}$ denote the number of true positives (overlapping tokens between the predicted and ground-truth answers), false positives (tokens predicted but not in the ground truth), and false negatives (tokens in the ground truth that were not predicted), respectively.
$F_{1}=2\cdot\frac{\mathrm{Precision}\cdot\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}.$
$F_{1}=\frac{2\cdot\mathrm{TP}}{2\cdot\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}.$
Baselines. We primarily compare with the following baselines from wu2023defending; hines2024defending; schulhoff2023ignore. Vanilla: Original prompt without any defense technique. Delimiting: Adding special characters at the start and end of the context. Datamarking: Replace every space in the context with a special character. Sandwich: Wrap the context with user instruction as a sandwich. Encode_\_Base64: The context is encoded into Base64 while the other text spans are provided with plain text. For fair comparison, we do not compare with baselines of finetuning or requiring test-time computation with extra LLM calls per response.
Our CachePrune is implemented based on Vanilla. We should note that our CachePrune is actually complementary to the other baselines, since our approach does not modify the prompt. In addition, its is unfair comparing CachePrune with other approaches based on fine-tuning or test-time workflows (Section 3). This is because such approaches require much larger computation cost than ours. Specifically, fine-tuning requires at least hundreds of samples while we only need < 10 samples. Our approach only requires a single forward pass for each testing sample while the test-time workflows require multiple LLM calls.
Appendix B Insights from DPO
In this section, we want to derive a deeper intuition on our preferential attribution loss $\mathcal{L}^{attr}_{full}$, by analyzing the DPO loss $\mathcal{L}_{DPO}$. Specifically, we first present an upperbound of $\mathcal{L}_{DPO}$ in the context of indirect prompt injection attack (Theorem 1). This upperbound helps us understand the objective of preference optimization in defending against such attack, i.e., by decomposing the preference objective into the terms of Probability and Uniformity. We show that our attribution loss is closely associated with these two terms. Further, to validate such an association, we show in Lemma 1 that our attribution loss is consistent with the DPO upperbound in the asymptotic cases. Specifically, sthe upperbound is getting positive infinite when our attribution loss is taking its largest value (i.e., 1), and vise verse.
where $|\cdot|$ is the support of a distribution. $\mathcal{C}_{ref,\mathcal{D}}$ is a constant to $\theta$ that is functioned by $\mathcal{D}$ and the DPO reference model refref. $\mathbb{H}(\mathcal{Y}^{c}_{x}|x)$ and $\mathbb{H}(\mathcal{Y}^{p}_{x}|x)$ are the entropy of clean and poisoned responses given $x$. $p_{\theta}(y\in|\mathcal{Y}^{p}_{x}|\,|x)$ is the gross probability of generating poisoned responses from $x$, and similar to $p_{\theta}(y\in|\mathcal{Y}^{c}_{x}|\,|x)$. The proof is in Appendix C.
(a) LLama3-8B
(b) Mistral-7B
ASR $\downarrow$ F1 (clean) $\uparrow$ F1 (attack) $\uparrow$
k=1 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
k=2 5.57 $\pm$ 0.30 26.03 $\pm$ 0.28 22.47 $\pm$ 0.37
k=4 10.77 $\pm$ 0.45 24.71 $\pm$ 0.37 19.29 $\pm$ 0.33
$\mathcal{L}^{attr}_{full}$ 14.81 $\pm$ 0.59 25.63 $\pm$ 0.39 19.78 $\pm$ 0.55
ASR $\downarrow$ F1 (clean) $\uparrow$ F1 (attack) $\uparrow$
$\alpha$=1.5 6.40 $\pm$ 0.32 26.71 $\pm$ 0.53 20.22 $\pm$ 0.56
$\alpha$=1.0 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
$\alpha$=0.5 10.77 $\pm$ 0.61 28.33 $\pm$ 0.40 21.29 $\pm$ 0.61
$\alpha$=0.3 13.50 $\pm$ 0.70 28.91 $\pm$ 0.37 21.78 $\pm$ 0.43
Especially, the clean and poison probabilities $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$ are not sufficient to capture the objective of preference optimization. In order to minimize (14), we should also attend to the uniformity with $\mathbb{H}(\mathcal{Y}^{p/c}_{x}|x)$. This requires computing the expectation over the generated responses,which is sample inefficient due to the complexity of the space of generated responses. Here, we delegate the entropy terms with the most probable poison and clean responses, denoted as $y^{p/c,*}_{x}=\text{argmax}_{y\in|\mathcal{Y}_{x}^{p/c}|}p_{\theta}(y|x)$. Intuitively, given the gross probability $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$, $\mathbb{H}(\mathcal{Y}^{p/c}_{x}|x)$ should be generally lowered if $y^{p/c,*}_{x}$ gets higher probability, vice versa.
We can observe that the $\mathcal{L}_{full}^{attr}$ (8) captures both the objectives of probability and uniformity in (14): a) Minimizing (8) promotes $p_{\theta}(y\in|\mathcal{Y}^{c}_{x}|\,|x)$, while suppressing $p_{\theta}(y\in|\mathcal{Y}^{p}_{x}|\,|x)$. b) Given the gross probability $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$, we have
Lemma 1. As $\mathcal{L}^{attr}_{full}$ is ranged between [1,1][-1,1], it is closely associated with $\mathcal{L}_{DPO}^{u}$ by,
We can observe from the proof that Lemma 1 will no longer hold with only few samples, if we follow (7) that replace $y^{p/c,*}$ with $y^{p/c}$ in $\mathcal{L}_{full}^{attr}$. This suggests to sample with the most probable responses for feature attribution.
Appendix C Proof of Theorem 1
w/ $\Phi$ ASR $\downarrow$ F1 (clean) $\uparrow$ F1 (attack) $\uparrow$
pp=0.5% Y 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
N 7.65 $\pm$ 0.63 29.06 $\pm$0.42 21.35 $\pm$ 0.44
pp=1.0% Y 4.38 $\pm$ 0.32 28.12 $\pm$ 0.58 23.17 $\pm$ 0.30
N 4.64 $\pm$ 0.53 26.01 $\pm$ 0.43 18.77 $\pm$ 0.91
pp=5.0% Y 0.83 $\pm$ 0.07 27.48 $\pm$ 0.66 23.03 $\pm$ 0.64
N 7.86 $\pm$ 0.86 6.48 $\pm$ 1.04 10.43 $\pm$ 2.05
Table 5: Ablation on whether to prune from $\Phi$ using SQuAD with LLaMA-3-8B. “Y” means that (5) and (6) only mask and prune the neurons defined in (LABEL:eq:_phi). “N” means that $\Phi$ represents all neurons in the KV cache. Pruning from $\Phi$ becomes increasingly important as more neurons are pruned ($p\uparrow$), i.e., when robustness requirements become stricter relative to utility.
NN ASR $\downarrow$ F1 (clean) $\uparrow$ F1 (attack) $\uparrow$
Llama3-8B 44 8.15 $\pm$ 0.58 26.14 $\pm$ 0.76 22.99 $\pm$ 1.19
88 7.44 $\pm$ 0.22 28.68 $\pm$ 0.30 22.84 $\pm$ 0.49
1212 7.52 $\pm$ 0.31 28.27 $\pm$ 0.42 24.12 $\pm$ 0.50
Mistral-7B 44 0.74 $\pm$ 0.36 24.75 $\pm$ 0.65 22.52 $\pm$ 1.67
88 0.68 $\pm$ 0.41 24.46 $\pm$ 0.91 23.10 $\pm$ 1.32
1212 0.57 $\pm$ 0.32 25.05 $\pm$ 1.05 23.26 $\pm$ 0.88
Phi-3.5-mini- instruct (3.8B) 44 0.86 $\pm$ 0.32 25.76 $\pm$ 0.77 27.12 $\pm$ 1.31
88 0.71 $\pm$ 0.18 26.76 $\pm$ 0.56 25.55 $\pm$ 0.60
1212 0.62 $\pm$ 0.13 26.37 $\pm$ 0.71 25.26 $\pm$ 0.47
Table 6: Performance with number of samples NN used for attribution. We observe that the ASR generally decreases as NN increases, with the variance also showing a decreasing trend.